40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1709     Entry time: Tue Jun 30 23:09:40 2009
Author: Alberto 
Type: Update 
Category: Locking 
Subject: chronicles of some locking attempts 

Tonight I tried to lock the interferometer. At the first attempts the arm power didn't go above about 4. The mode cleaner seemed to be not well aligned and it lost lock or got stuck on a wrong mode. I had to run the MC_UP and MC_DOWN scripts to lock it again.

After that the locking proceed more smoothly; at least till a power level in the arms of about 60. Then again the mode cleaner lost lock and I had to run the scripts again. Without the MCWFS servo off the MC reflected power is still rather high (about 1.7); also even when the WFS servo is engaged the reflected power is about 0.5, versus 0.3 that it should be.

Those are both signs of a not very good alignment. Tomorrow I'll have to work on the injection periscope on the PSL table to try to fix that.

ELOG V3.1.3-